OpenStudy (anonymous):

Please help i dont get this at all im really confused :P 1. Dolora spends at least 4 hours per week gardening. she spends more time on routine maintenance such as weeding, pruing , and watering than on harvasing vegtables or flowers . she spends at least 1 hour per week harvesting. which system of inequalities represents Dolora's weekly gardening time if x represents the number of weekly hours spend on routine garden maintenance and y represents the number of weekly hours spent on harvesting?

OpenStudy (anonymous):

A. \[\left\{ x+y >4 \right\} \] \[\left\{ x >y \right\}\] \[\left\{ y \ge1 \right\}\]

OpenStudy (anonymous):

This was one of the hardest one on the test. it's x+y>=4 x>y y>=1 It's c.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You're welcome. Good luck on the test. It took me forever to get done. :)

OpenStudy (anonymous):

lol its taking me a long time im barley one #12

OpenStudy (anonymous):

lol. They're killing us with these super long tests.

OpenStudy (anonymous):

i know right :P